# Simple is Effective: The Roles of Graphs and Large Language Models in Knowledge-Graph-Based Retrieval-Augmented Generation

### v2

@@ Abstract
Large Language Models (LLMs) demonstrate strong reasoning abilities but face limitations such as hallucinations and outdated knowledge. Knowledge Graph (KG)-based Retrieval-Augmented Generation (RAG) addresses these issues by grounding LLM outputs in structured external knowledge from KGs. However, current KG-based RAG frameworks still struggle to optimize the trade-off between retrieval effectiveness and efficiency in identifying a suitable amount of relevant graph information for the LLM to digest. We introduce SubgraphRAG, extending the KG-based RAG framework that retrieves subgraphs and leverages LLMs for reasoning and answer prediction. Our approach innovatively integrates a lightweight multilayer perceptron with a parallel triple-scoring mechanism for efficient and flexible subgraph retrieval while encoding directional structural distances to enhance retrieval effectiveness. The size of retrieved subgraphs can be flexibly adjusted to match the query’s need and the downstream LLM’s capabilities. This design strikes a balance between model complexity and reasoning power, enabling scalable and generalizable retrieval processes. Notably, based on our retrieved subgraphs, smaller LLMs like Llama3.1-8B-Instruct deliver competitive results with explainable reasoning, while larger models like GPT-4o achieve state-of-the-art accuracy compared with previous baselines—all without fine-tuning. Extensive evaluations on the WebQSP and CWQ benchmarks highlight SubgraphRAG’s strengths in efficiency, accuracy, and reliability by reducing hallucinations and improving response grounding. Our implementation is available at https://github.com/Graph-COM/SubgraphRAG.
@@ 1 Introduction
Large language models (LLMs) have increasingly demonstrated remarkable reasoning capabilities across various domains (Brown et al., 2020; Kojima et al., 2022; Wei et al., 2022; Bubeck et al., 2023; Yao et al., 2023; Huang & Chang, 2023). However, issues like hallucinations (Ji et al., 2023; Huang et al., 2023; Zhang et al., 2023), outdated knowledge (Dhingra et al., 2022; Kasai et al., 2023), and a lack of vertical, domain-specific expertise (Li et al., 2023b) undermine the trustworthiness of LLM outputs. Retrieval-augmented generation (RAG) has emerged as a promising strategy to mitigate these problems by grounding LLM outputs in external knowledge sources (Shuster et al., 2021; Borgeaud et al., 2022; Vu et al., 2024; Gao et al., 2024b).
@@
Despite the effectiveness of text-based retrieval, graph structures offer a more efficient alternative for organizing knowledge (Chein & Mugnier, 2008). Graphs facilitate explicit representation of relationships, reduce information redundancy, and allow for more flexible updates (Robinson et al., 2015). Recent studies have explored using graph-structured knowledge, particularly knowledge graphs (KGs), as external resources for RAG (Pan et al., 2024; Peng et al., 2024; Edge et al., 2024). However, developing effective and efficient frameworks for KG-based RAG remains limited due to the unique challenges involved in retrieving information from the complex structures of KGs.
@@
Firstly, traditional text-based retrieval methods, such as BM25 (Robertson et al., 1994; Robertson & Zaragoza, 2009) or dense retrieval with cosine similarity (Karpukhin et al., 2020), are insufficient for supporting LLMs in complex reasoning tasks (Sun et al., 2018). For instance, a query like “What is the most famous painting by a contemporary of Michelangelo and Raphael?” requires not only retrieving works by their contemporaries but also reasoning about relationships beyond Michelangelo and Raphael themselves. Thus, KG retrieval goes beyond basic entity linking, requiring the extraction of nonlocal entities connected through multi-hop, relevant relationships to support reasoning (Jiang et al., 2023a; Luo et al., 2024; Sun et al., 2024a). Such information is often best represented as KG subgraphs, whose retrieval enables more effective downstream reasoning.
@@
Figure 1: The framework of SubgraphRAG. Retrieved subgraphs consist of relevant triples that are extracted in parallel. Retrieved subgraphs are flexible in their forms and their sizes. In the above example, the relevant subgraph has flexible and complex forms (neither trees nor paths).
@@
Second, KG-based RAG faces significant computational challenges. Traditional efficient search methods, such as locality-sensitive hashing, which are designed for similarity search, are not well-suited for extracting complex structural patterns such as paths or subgraphs. With the need to handle potential online graph queries and adapt to dynamic updates in KGs (Trivedi et al., 2017; Liang et al., 2024), efficiently identifying relevant structural information while meeting latency requirements is crucial for designing a practical KG-based RAG framework.
@@
Third, extracting structural information must cover the critical evidence needed to answer the query without exceeding the reasoning capacity of LLMs. Expanding the context window increases computational complexity and can degrade RAG performance by introducing irrelevant information (Xu et al., 2024) and causing the “lost in the middle” phenomenon (Liu et al., 2024b). To prevent these issues, redundant structural information should be pruned to keep only relevant evidence within the LLMs’ processing limits, improving accuracy and avoiding hallucinations of LLMs.
@@
Existing KG-based RAG frameworks face limitations in addressing the aforementioned challenges, often due to suboptimal balancing between information retrieval and reasoning over the retrieved data. For example, many approaches rely on LLMs to perform retrieval through step-by-step searches from entities to their neighbors over KGs, resulting in significant complexity by requiring multiple LLM calls (e.g., GPT-4) for each query (Kim et al., ; Gao et al., 2024a; Wang et al., 2024; Guo et al., 2024; Ma et al., 2024; Sun et al., 2024a; Jiang et al., 2024; Jin et al., 2024). These methods may also miss relevant entities or relationships due to the vast search space over KGs and the limited context windows of LLMs. Conversely, methods that employ lighter models for retrieval, such as LSTMs or GNNs, embed iterative reasoning within the retrieval process itself (Zhang et al., 2022; Liu et al., 2024a; Sun et al., 2019). While more efficient, these methods are constrained by the limited reasoning capacity of lighter models, which can lead to the omission of crucial evidence needed to answer queries. Additionally, some approaches retrieve fixed types of subgraphs for efficiency such as paths (Zhang et al., 2022; Luo et al., 2024), but this restricts the coverage of critical evidence needed for LLM reasoning—a point we will explore more in Sec. 3.1.
@@
Design Principles We argue that there is an inherent tradeoff between model complexity and reasoning capability. To effectively search over KGs, which are expected to grow rapidly, knowledge retrievers should remain lightweight, flexible, generalizable, and equipped with basic reasoning abilities to efficiently filter relevant (even if only roughly relevant) information from vast amounts of irrelevant data, while delegating complex reasoning tasks to LLMs. As LLMs continue to demonstrate increasingly sophisticated reasoning capabilities and are likely to improve further, this division of labor becomes more reasonable. As long as the retrieved information fits within the LLM’s reasoning capacity, LLMs—leveraging their superior reasoning power—can then perform more fine-grained analysis and provide accurate answers with appropriate prompting. This approach extends the two-stage Recall & Ranking (rough to fine) framework commonly used in the traditional pipelines of information retrieval and recommendation, with the key advance being that each stage is paired with appropriately tiered reasoning capabilities from AI models to meet the demands of responding to complex queries. Besides, for questions proven too challenging, the concept of iterating the above process can be adopted. However, this consideration is beyond the scope of the current work.
@@
Present Work Our KG-based RAG framework, SubgraphRAG (Fig. 1), follows a pipeline that first retrieves a relevant subgraph and then employs LLMs to reason over it. While this approach mirrors some existing methods (He et al., 2021; Jiang et al., 2023b), SubgraphRAG introduces novel design elements that significantly improve both efficiency and effectiveness by adhering to the aforementioned principles. For efficiency, we employ a lightweight multilayer perceptron (MLP) combined with parallel triple-scoring for subgraph retrieval. To ensure effectiveness, we encode tailored structural distances from the topic entities of a query as structural features. This enables our MLP retriever to outperform more complex models, such as GNNs, LLMs, and heuristic searches, in terms of covering the triples and entities critical for answering the query while maintaining high efficiency. Additionally, the retrieved subgraphs have flexible forms, with adjustable sizes to accommodate the varying capacities of LLMs. SubgraphRAG employs unfine-tuned LLMs, maintaining generalization, adaptability to updated KGs, and compatibility with black-box LLMs.
@@
We evaluate SubgraphRAG on two prominent multi-hop knowledge graph question answering (KGQA) benchmarks—WebQSP and CWQ. Remarkably, without fine-tuning, smaller models like Llama3.1-8B-Instruct can achieve competitive performance. Larger models, such as GPT-4o, deliver state-of-the-art (SOTA) results, surpassing previous methods for most cases. Furthermore, SubgraphRAG shows robust multi-hop reasoning capabilities, excelling on more complex, multi-hop questions and demonstrating effective generalization across datasets despite domain shifts. Ablation studies on different retrievers highlight the advantage of our retriever, which consistently outperforms baseline retrievers and is key to our superior KGQA performance. Additionally, our method also exhibits a substantial capability of reducing hallucination by generating knowledge-grounded answers and explanations for its reasoning.
@@ 2 Preliminaries
KG-based RAG aims to enhance LLM responses by incorporating knowledge from a KG as contextual information. Given a query $q$, LLMs can access relevant knowledge represented by triples in the KG to address the request posed by $q$. The challenge lies in efficiently searching for relevant knowledge within the often large-scale KG and reasoning to generate an accurate response.
Entity Linking Entity linking is often the first step in KG-based RAG, whose goal is to identify the set of entities ${\mathcal{T}}_{q}\subset{\mathcal{E}}$ directly involved in the query $q$. The entities in ${\mathcal{T}}_{q}$, named topic entities, provide valuable inductive bias for retrieval as the triples relevant to $q$ are often close to ${\mathcal{T}}_{q}$.
Knowledge Graph Question Answering (KGQA) is a key application often used to evaluate KG-based RAG, where the query $q$ is a question that requires finding answers under specific constraints. The answer(s) ${\mathcal{A}}_{q}$ typically corresponds to a set of entities in the KG. Questions that require multiple triples in the KG as evidence to identify an answer entity are classified as complex questions, as they demand multi-hop reasoning, in contrast to single-hop questions.
@@ 3 The SubgraphRAG framework
Our proposed framework, SubgraphRAG, adopts a retrieval-and-reasoning pipeline for KG-based RAG. Specifically, given a query $q$, SubgraphRAG first extracts a subgraph ${\mathcal{G}}_{q}\subset{\mathcal{G}}$ that represents relevant knowledge, and then LLMs generate the response by reasoning over ${\mathcal{G}}_{q}$. While some existing works follow a similar pipeline (Zhang et al., 2022; Luo et al., 2024; Wu et al., 2023; Wen et al., 2023), our approach is novel in addressing three key challenges: First, the extracted subgraph ${\mathcal{G}}_{q}$ is designed to cover as much of the relevant evidence for answering $q$ as possible, while adhering to a size constraint $K$, which can be adjusted according to the capacity of the downstream LLM. Second, the extraction of ${\mathcal{G}}_{q}$ is highly efficient and scalable. Third, we employ tailored prompting to guide the LLM in reasoning over ${\mathcal{G}}_{q}$ and generating a well-grounded answer with explanations.
@@ 3.1 Efficient, Flexible and Expressive Subgraph Retrieval
Problem Reduction To begin with, we formulate the subgraph retrieval problem and gradually reduce it to an efficiently solvable problem. An LLM can be viewed as an answer generator $\mathbb{P}(\cdot|{\mathcal{G}}_{q},q)$ that takes queries and evidence represented by subgraphs. Given a query $q$ and its answer $\mathcal{A}_{q}$, the best subgraph evidence for this LLM is denoted as ${\mathcal{G}}_{q}^{*}=\operatorname*{arg\,max}_{{\mathcal{G}}_{q}\subseteq{\mathcal{G}}}\mathbb{P}(\mathcal{A}_{q}\mid{\mathcal{G}}_{q},q)$. Of course, solving this problem is practically impossible as it requires the knowledge of $\mathcal{A}_{q}$. Instead, we aim to learn a subgraph retriever from data and expect this retriever to generalize to unseen future queries.
Specifically, let the subgraph retriever be a distribution $\mathbb{Q}_{\theta}(\cdot|q,{\mathcal{G}})$ over the subgraph space of the KG. $\theta$ denotes the parameters. Given a training set of question-answer pairs $\mathcal{D}$, the subgraph retriever learning problem can be formulated as the following problem:
@@
In practice, this problem is still hard to solve due to the complexity of the LLM, i.e., the form of $\mathbb{P}$. Simply evaluating $\mathbb{P}(\mathcal{A}_{q}\mid{\mathcal{G}}_{q},q)$ means calling the LLM to generate the particular answer $\mathcal{A}_{q}$, which could be costly and only applicable to grey/white-box LLMs with accessible output logits, let alone the incomputable gradient $\frac{d\mathbb{P}}{d{\mathcal{G}}_{q}}$.
To solve the problem in Eq. 1, we adopt the following idea. If we know the optimal subgraph ${\mathcal{G}}_{q}^{*}$, the maximum likelihood estimation (MLE) principle can be leveraged to train the retriever $\max_{\theta}\;\mathbb{E}_{(q,\mathcal{A}_{q})\sim\mathcal{D}}\;\mathbb{Q}_{\theta}({\mathcal{G}}_{q}^{*}\mid{\mathcal{G}},q)$. However, getting ${\mathcal{G}}_{q}^{*}$ for an even known question-answer pair $(q,\mathcal{A}_{q})$ is computationally hard and LLM-dependent. Instead, we use $(q,\mathcal{A}_{q})$ to construct surrogate subgraph evidence with heuristics $\tilde{{\mathcal{G}}}(q,\mathcal{A}_{q})$ and train the retriever based on MLE:
@@
Some examples of $\tilde{{\mathcal{G}}}_{q}$ could be the shortest paths between topic entities ${\mathcal{T}}_{q}$ and the answer entities ${\mathcal{A}}_{q}$. Eq. 2 is conceptually similar to the weak supervision adopted in some existing work (Zhang et al., 2022). However, the formulation in Eq. 2 indicates that the sampled subgraph does not necessarily follow a fixed type (trees or paths). Instead, the retriever distribution $\mathbb{Q}_{\theta}$ can by construction factorize into a product of distributions over triples, allowing efficient training and inference, flexible subgraph forms, and adjustable subgraph sizes.
@@
$\mathbb{Q}_{\theta}({\mathcal{G}}_{q}\mid{\mathcal{G}},q)=\prod_{\tau\in{\mathcal{G}}_{q}}p_{\theta}(\tau\mid z_{\tau},q)\prod_{\tau\in{\mathcal{G}}\backslash{\mathcal{G}}_{q}}(1-p_{\theta}(\tau\mid z_{\tau},q)),$
@@
Relevant Designs Previous approaches often adopt heuristics and focus on some particular types of subgraphs, such as constrained subgraph search (e.g., searching for connected subgraphs (He et al., 2024)), constrained path search from topic entities, often imposing constraints in path counts and lengths and employing expensive iterative processes (Zhang et al., 2022; Wu et al., 2023; Luo et al., 2024; Sun et al., 2024a; Liu et al., 2024a; Mavromatis & Karypis, 2024; Sun et al., 2024b), and entity selection followed by extracting entity-induced subgraphs, where all triples involving an entity are included together (Yasunaga et al., 2021; Taunk et al., 2023). The loss in flexibility narrows the space of possible retrieved subgraphs, which eventually harms the effectiveness of the RAG.
@@
Directional Distance Encoding (DDE) as $z_{\tau}({\mathcal{G}},q)$ The latent variable $z_{\tau}({\mathcal{G}},q)$ aims to model the relationship between a triple $\tau$ and the query $q$ given ${\mathcal{G}}$. One idea is to employ graph neural networks (GNNs) to compute $z_{\tau}({\mathcal{G}},q)$ through message passing between entities/relations with attribute embeddings and question embeddings. Some previous works indeed adopt GNNs to get latent representations of entities (Yasunaga et al., 2021; Kang et al., 2023; Mavromatis & Karypis, 2024; Liu et al., 2024a). However, GNNs are known to have limited representation power (Xu et al., 2019; Morris et al., 2019; Chen et al., 2020).
@@
The structural relationship between $\tau$ and $q$ provides valuable information complementing to their semantic relationship. Inspired by the success of distance encoding and labeling trick in enhancing the structural representation power of GNNs (Li et al., 2020; Zhang et al., 2021), we propose a DDE as $z_{\tau}({\mathcal{G}},q)$ to model the structural relationship. Given topic entities ${\mathcal{T}}_{q}$, let $\textbf{s}_{e}^{(0)}$ be a one-hot encoding representing $e\in{\mathcal{T}}_{q}$ or $e\not\in{\mathcal{T}}_{q}$. For the $l+1$-th round, we perform feature propagation and compute $\textbf{s}_{e}^{(l+1)}=\text{MEAN}\{\textbf{s}_{e^{\prime}}^{(l)}\mid(e^{\prime},\cdot,e)\in{\mathcal{G}}\}$, and through the reverse direction to account the directed nature of ${\mathcal{G}}$, $\textbf{s}_{e}^{(r,l+1)}=\text{MEAN}\{\textbf{s}_{e^{\prime}}^{(r,l)}\mid(e,\cdot,e^{\prime})\in{\mathcal{G}}\}$, where $\textbf{s}_{e}^{(r,0)}=\textbf{s}_{e}^{(0)}$. We concatenate the results across all rounds and both directions to obtain the final entity encodings $\textbf{s}_{e}=[\textbf{s}_{e}^{(0)}\|\textbf{s}_{e}^{(1)}\|\cdots\|\textbf{s}_{e}^{(r,1)}\|\cdots]$, which leads to triple encodings as $z_{\tau}({\mathcal{G}},q)=[\textbf{s}_{h}||\textbf{s}_{t}]$ that concatenates the head $h$’s and the tail $t$’s encodings. In section 4.1, we compare different approaches to compute $z_{\tau}({\mathcal{G}},q)$ using GNNs, DDEs or only one-hot encodings of ${\mathcal{T}}_{q}$, and DDEs perform the best.

### v3

@@ Abstract
Large Language Models (LLMs) demonstrate strong reasoning abilities but face limitations such as hallucinations and outdated knowledge. Knowledge Graph (KG)-based Retrieval-Augmented Generation (RAG) addresses these issues by grounding LLM outputs in structured external knowledge from KGs. However, current KG-based RAG frameworks still struggle to optimize the trade-off between retrieval effectiveness and efficiency in identifying a suitable amount of relevant graph information for the LLM to digest. We introduce SubgraphRAG, extending the KG-based RAG framework that retrieves subgraphs and leverages LLMs for reasoning and answer prediction. Our approach innovatively integrates a lightweight multilayer perceptron (MLP) with a parallel triple-scoring mechanism for efficient and flexible subgraph retrieval while encoding directional structural distances to enhance retrieval effectiveness. The size of retrieved subgraphs can be flexibly adjusted to match the query’s needs and the downstream LLM’s capabilities. This design strikes a balance between model complexity and reasoning power, enabling scalable and generalizable retrieval processes. Notably, based on our retrieved subgraphs, smaller LLMs like Llama3.1-8B-Instruct deliver competitive results with explainable reasoning, while larger models like GPT-4o achieve state-of-the-art accuracy compared with previous baselines—all without fine-tuning. Extensive evaluations on the WebQSP and CWQ benchmarks highlight SubgraphRAG’s strengths in efficiency, accuracy, and reliability by reducing hallucinations and improving response grounding. Our implementation is available at https://github.com/Graph-COM/SubgraphRAG.
@@ 1 Introduction
Large language models (LLMs) have increasingly demonstrated remarkable capabilities (Brown et al., 2020; Kojima et al., 2022; Wei et al., 2022; Bubeck et al., 2023; Yao et al., 2023; Huang & Chang, 2023). However, issues like hallucinations (Ji et al., 2023; Huang et al., 2023; Zhang et al., 2023), outdated knowledge (Dhingra et al., 2022; Kasai et al., 2023), and a lack of domain expertise (Li et al., 2023b) undermine their trustworthiness. Retrieval-augmented generation (RAG) has emerged as a promising strategy to mitigate these problems by grounding LLM outputs in external knowledge sources (Shuster et al., 2021; Borgeaud et al., 2022; Vu et al., 2024; Gao et al., 2024b).
@@
Despite the effectiveness of text-based RAG, graph-based structures offer a more efficient alternative for organizing knowledge (Chein & Mugnier, 2008). Graphs facilitate explicit representation of relationships, reduce information redundancy, and allow for more flexible updates (Robinson et al., 2015). Recent studies have explored using graph-structured knowledge, particularly knowledge graphs (KGs), as RAG resources (Pan et al., 2024; Peng et al., 2024; Edge et al., 2024). However, developing effective and efficient frameworks for KG-based RAG remains limited due to the unique challenges posed by information retrieval from KGs.
@@
Firstly, traditional text retrieval methods, such as BM25 (Robertson et al., 1994; Robertson & Zaragoza, 2009) and dense retrieval (Karpukhin et al., 2020), are insufficient for supporting LLMs in complex reasoning tasks (Sun et al., 2018). For instance, a query like “What is the most famous painting by a contemporary of Michelangelo and Raphael?” requires not only retrieving works by their contemporaries but also reasoning about relationships beyond Michelangelo and Raphael themselves. Thus, KG retrieval goes beyond basic entity linking, requiring the extraction of nonlocal entities connected through multi-hop and semantically relevant relationships to support reasoning (Jiang et al., 2023a; Luo et al., 2024b; Sun et al., 2024a). Such information is often best represented as KG subgraphs, whose retrieval enables more effective downstream reasoning.
@@
Figure 1: The SubgraphRAG framework. Retrieved subgraphs are flexible in form and size, consisting of triples extracted in parallel. This example subgraph has a complex form beyond a path.
@@
Second, KG-based RAG faces significant computational challenges. In order to handle online queries and adapt to dynamic KG updates (Trivedi et al., 2017; Liang et al., 2024), efficiently identifying relevant structural information while meeting latency requirements is crucial for designing a practical KG-based RAG framework. Efficient approximate methods for nearest neighbor search like locality-sensitive hashing (Indyk & Motwani, 1998) are not suitable for structure extraction.
@@
Third, retrieved structural information must cover the critical evidence needed to answer the query while respecting the reasoning capacity of LLMs. Simply expanding the context window increases computational cost and can degrade RAG performance due to the introduction of irrelevant information (Xu et al., 2024), causing the “lost in the middle” phenomenon (Liu et al., 2024b). To prevent these issues, redundant structural information should be pruned to keep only relevant evidence within the LLMs’ processing limits, improving accuracy and avoiding hallucinations of LLMs.
@@
Existing KG-based RAG frameworks often struggle to effectively balance retrieval with reasoning over the retrieved data, limiting their ability to address these challenges. For example, many approaches rely on multiple LLM calls for iterative KG search, introducing significant complexity for each query (Kim et al., ; Gao et al., 2024a; Wang et al., 2024; Guo et al., 2024; Ma et al., 2024; Sun et al., 2024a; Jiang et al., 2024; Jin et al., 2024; Xiong et al., 2024). Furthermore, the mismatch between large KG search spaces and constrained LLM context windows poses challenges to the coverage of LLM-based retrieval. Conversely, some methods employ lighter models, such as Long Short-Term Memory or Graph Neural Networks, for both retrieval and reasoning (Zhang et al., 2022; Liu et al., 2024a; Sun et al., 2019). Despite being more efficient, these methods are constrained by the inferior reasoning capabilities of the lighter models. Additionally, certain approaches retrieve fixed types of subgraphs (e.g., paths) for convenience (Zhang et al., 2022; Luo et al., 2024b), which restricts the coverage of critical evidence for LLM reasoning—a point we explore further in Sec. 3.1.
@@
Design Principles We argue that there is an inherent tradeoff between model complexity and reasoning capability. To effectively search over KGs, which are expected to grow rapidly, knowledge retrievers should remain lightweight, flexible, generalizable, and equipped with basic reasoning abilities to efficiently filter (roughly) relevant information from vast amounts of irrelevant data while delegating complex reasoning tasks to LLMs. As LLMs continue to demonstrate increasingly sophisticated reasoning capabilities and are likely to improve further, this division of labor will become more sensible. As long as the retrieved information fits within the LLM’s reasoning capacity, LLMs—leveraging their superior reasoning power—can then perform more fine-grained analysis and provide accurate answers with appropriate prompting. This approach extends the two-stage Recall & Ranking (rough to fine) framework commonly used in traditional information retrieval and recommendation pipelines, with the key advance being that each stage is paired with appropriately tiered reasoning capabilities from AI models to meet the demands of answering complex queries. Besides, for questions proven too challenging, the concept of iterating the above process can be adopted. However, this consideration is beyond the scope of the current work.
@@
Present Work Our KG-based RAG framework, SubgraphRAG (Fig. 1), follows a pipeline that first retrieves a relevant subgraph and then employs LLMs to reason over it. While this approach mirrors some existing methods (He et al., 2021; Jiang et al., 2023b), SubgraphRAG introduces novel design elements that significantly improve both efficiency and effectiveness by adhering to the aforementioned principles. For efficiency, we employ a lightweight multilayer perceptron (MLP) combined with parallel triple-scoring for subgraph retrieval. To ensure effectiveness, we encode tailored structural distances from the query topic entities as structural features. This enables our MLP retriever to outperform more complex models, such as GNNs, LLMs, and heuristic searches, in terms of covering the triples and entities critical for answering the query while maintaining high efficiency. Additionally, the retrieved subgraphs have flexible forms, with adjustable sizes to accommodate the varying capacities of LLMs. SubgraphRAG employs unfine-tuned LLMs, maintaining generalization, adaptability to updated KGs, and compatibility with black-box LLMs.
@@
We evaluate SubgraphRAG on two multi-hop knowledge graph question answering benchmarks—WebQSP and CWQ. Remarkably, without fine-tuning, smaller models like Llama3.1-8B-Instruct can achieve competitive performance. Larger models, such as GPT-4o, deliver state-of-the-art results. Furthermore, SubgraphRAG shows robust multi-hop reasoning capabilities, excelling on more complex, multi-hop questions and demonstrating effective generalization across datasets despite domain shifts. Ablation studies on different retrievers highlight the advantage of our retriever, which consistently outperforms baseline retrievers and is key to our superior KGQA performance. Additionally, our method also exhibits a substantial capability of reducing hallucination by generating knowledge-grounded answers and explanations for its reasoning.
@@ 2 Preliminaries
KG-based RAG enhances LLMs with KG knowledge. Given a query $q$, LLMs can access relevant knowledge represented by triples in the KG to address $q$. The challenge lies in efficiently searching for relevant knowledge within the often large-scale KG and reasoning for accurate prediction.
Entity Linking is often the first step in KG-based RAG, whose goal is to identify the set of entities ${\mathcal{T}}_{q}\subset{\mathcal{E}}$ directly involved in the query $q$. The entities in ${\mathcal{T}}_{q}$, named topic entities, provide valuable inductive bias for retrieval as the triples relevant to $q$ are often close to ${\mathcal{T}}_{q}$.
Knowledge Graph Question Answering (KGQA) is an important application often used to evaluate KG-based RAG, where the query $q$ is a question that requires finding answers that satisfy certain constraints. The answer(s) ${\mathcal{A}}_{q}$ often correspond to a set of entities in the KG. Questions that require multiple triples in the KG as evidence to identify an answer entity are classified as complex questions, as they demand multi-hop reasoning, in contrast to single-hop questions.
@@ 3 The SubgraphRAG framework
Our proposed framework, SubgraphRAG, adopts a retrieval-and-reasoning pipeline for KG-based RAG. Specifically, given a query $q$, SubgraphRAG first extracts a subgraph ${\mathcal{G}}_{q}\subset{\mathcal{G}}$ that represents predicted relevant knowledge, and then LLMs generate the response by reasoning over ${\mathcal{G}}_{q}$. While some existing works adopt a similar pipeline (Zhang et al., 2022; Luo et al., 2024b; Wu et al., 2023; Wen et al., 2023), our approach is novel in addressing three key challenges: First, the extracted subgraph ${\mathcal{G}}_{q}$ is designed to comprehensively cover the relevant evidence for answering $q$, while adhering to a size constraint $K$, which can be adjusted according to the capacity of the downstream LLM. Second, the extraction of ${\mathcal{G}}_{q}$ is highly efficient and scalable. Third, we employ tailored prompting to guide the LLM in reasoning over ${\mathcal{G}}_{q}$ and generating a well-grounded answer with explanations.
@@ 3.1 Efficient, Flexible and Expressive Subgraph Retrieval
Problem Reduction We formulate the subgraph retrieval problem and reduce it to an efficiently solvable problem. An LLM models $\mathbb{P}(\cdot|{\mathcal{G}}_{q},q)$ that takes queries and evidence subgraphs. Given a query $q$ and its answer $\mathcal{A}_{q}$, the best subgraph evidence for this LLM is denoted as ${\mathcal{G}}_{q}^{*}=\operatorname*{arg\,max}_{{\mathcal{G}}_{q}\subseteq{\mathcal{G}}}\mathbb{P}(\mathcal{A}_{q}\mid{\mathcal{G}}_{q},q)$. However, solving this problem is impossible as it requires the knowledge of $\mathcal{A}_{q}$. Instead, we aim to learn a subgraph retriever that generalizes to unseen future queries.
Specifically, let the subgraph retriever be a distribution $\mathbb{Q}_{\theta}(\cdot|q,{\mathcal{G}})$ over the subgraph space of the KG, where $\theta$ denotes the parameters. Given a training set of question-answer pairs $\mathcal{D}$, the subgraph retriever learning problem can be formulated as the following problem:
@@
This problem is hard to solve due to the complexity of the LLM ($\mathbb{P}$). Evaluating $\mathbb{P}(\mathcal{A}_{q}\mid{\mathcal{G}}_{q},q)$ requires the LLM to generate the particular answer $\mathcal{A}_{q}$, which can be costly and is only applicable to grey/white-box LLMs with accessible output logits, let alone the non-computable gradient $\frac{d\mathbb{P}}{d{\mathcal{G}}_{q}}$.
To solve the problem in Eq. 1, we adopt the following strategy. If we know the optimal subgraph ${\mathcal{G}}_{q}^{*}$, the maximum likelihood estimation (MLE) principle can be leveraged to train the retriever $\max_{\theta}\;\mathbb{E}_{(q,\mathcal{A}_{q})\sim\mathcal{D}}\;\mathbb{Q}_{\theta}({\mathcal{G}}_{q}^{*}\mid{\mathcal{G}},q)$. However, getting ${\mathcal{G}}_{q}^{*}$ for an even known question-answer pair $(q,\mathcal{A}_{q})$ is computationally hard and LLM-dependent. Instead, we use $(q,\mathcal{A}_{q})$ to construct surrogate subgraph evidence with heuristics $\tilde{{\mathcal{G}}}(q,\mathcal{A}_{q})$ and train the retriever based on MLE:
@@
An example $\tilde{{\mathcal{G}}}_{q}$ is the shortest paths between the topic entities ${\mathcal{T}}_{q}$ and answer entities ${\mathcal{A}}_{q}$. Eq. 2 is conceptually similar to the weak supervision in Zhang et al. (2022). However, the formulation in Eq. 2 indicates that the sampled subgraph does not necessarily follow a fixed type like path. Instead, the retriever distribution $\mathbb{Q}_{\theta}$ can by construction factorize into a product of distributions over triples, allowing efficient training and inference, flexible subgraph forms, and adjustable subgraph sizes.
@@
$\mathbb{Q}_{\theta}({\mathcal{G}}_{q}\mid{\mathcal{G}},q)=\prod_{\tau\in{\mathcal{G}}_{q}}p_{\theta}(\tau\mid z_{\tau},q)\prod_{\tau\in{\mathcal{G}}\backslash{\mathcal{G}}_{q}}\left(1-p_{\theta}\left(\tau\mid z_{\tau},q\right)\right),$ (3)
@@
Relevant Designs Previous approaches often adopt heuristics and focus on particular subgraph types, such as constrained subgraph search (e.g., connected subgraphs (He et al., 2024)), constrained path search from topic entities, often imposing constraints in path counts and lengths (Luo et al., 2024b; Mavromatis & Karypis, 2024) or employing potentially expensive iterative processes (Zhang et al., 2022; Wu et al., 2023; Sun et al., 2024a; Liu et al., 2024a; Sun et al., 2024b), and entity selection followed by extracting entity-induced subgraphs, where all triples involving an entity are included together (Yasunaga et al., 2021; Taunk et al., 2023). The loss in flexibility narrows the space of possible retrieved subgraphs, which eventually harms the RAG effectiveness.
@@
Directional Distance Encoding (DDE) $z_{\tau}({\mathcal{G}},q)$ models the relationship between a triple $\tau$ and the query $q$ given ${\mathcal{G}}$. Graph neural network (GNN) is one option for this modeling purpose through message passing with entity, relation, and question embeddings (Yasunaga et al., 2021; Kang et al., 2023; Mavromatis & Karypis, 2024; Liu et al., 2024a). However, it faces fundamental limitations in its representation power (Xu et al., 2019; Murphy et al., 2019; Morris et al., 2019; Chen et al., 2020). One critical limitation is its inability to compute distances between entities and topic entities (Li et al., 2020), which can be particularly problematic for multi-hop question answering.
@@
The structural relationship between $\tau$ and $q$ provides valuable information complementing their semantic relationship. Inspired by the success of distance encoding and labeling trick in enhancing the structural representation power of GNNs (Li et al., 2020; Zhang et al., 2021), we propose a DDE as $z_{\tau}({\mathcal{G}},q)$ to model the structural relationship. Given topic entities ${\mathcal{T}}_{q}$, let $\textbf{s}_{e}^{(0)}$ be a one-hot encoding representing whether $e\in{\mathcal{T}}_{q}$ or not. For the $l+1$-th round, we perform feature propagation and compute $\textbf{s}_{e}^{(l+1)}=\text{MEAN}\{\textbf{s}_{e^{\prime}}^{(l)}\mid(e^{\prime},\cdot,e)\in{\mathcal{G}}\}$, and through the reverse direction to account for the directed nature of ${\mathcal{G}}$, $\textbf{s}_{e}^{(r,l+1)}=\text{MEAN}\{\textbf{s}_{e^{\prime}}^{(r,l)}\mid(e,\cdot,e^{\prime})\in{\mathcal{G}}\}$, where $\textbf{s}_{e}^{(r,0)}=\textbf{s}_{e}^{(0)}$. We concatenate all results to obtain the final entity encodings $\textbf{s}_{e}=[\textbf{s}_{e}^{(0)}\|\textbf{s}_{e}^{(1)}\|\cdots\|\textbf{s}_{e}^{(r,1)}\|\cdots]$, leading to triple encodings $z_{\tau}({\mathcal{G}},q)=[\textbf{s}_{h}||\textbf{s}_{t}]$ that concatenate the encodings of $h$ and $t$. In section 4.1, we compare different approaches to compute $z_{\tau}({\mathcal{G}},q)$ using GNNs, DDE, or only one-hot encodings of ${\mathcal{T}}_{q}$, and DDEs perform the best. See Appendix A for additional illustrations and discussions.